OpenStudy (anonymous):

given the following sequence, find the missing term: __, __, 33,__, -1

OpenStudy (solomonzelman):

I must know is it supposed to be an ARITHMETIC or GEOMETRIC sequence ?

OpenStudy (anonymous):

arthrimetic :3

OpenStudy (anonymous):

this is what i have so far: t5=a+d(n-1) -1=a+d (5-1) -1= a+d(4) -1= a+4d

OpenStudy (solomonzelman):

Well, between `33` and `-1` you know that \(\normalsize\color{blue}{ 33~+2d~=-1}\) where ` d ` is the difference in the sequence. So can you find the difference ?

OpenStudy (anonymous):

that's the part where i don't get...

OpenStudy (solomonzelman):

What you don't get? Do you get why I am getting \(\normalsize\color{black}{ 33+2d=-1}\) ?

OpenStudy (anonymous):

no, i don't know how to get the difference

OpenStudy (solomonzelman):

subtract 33 from both sides and divide both sides by 2 . After you do this, d=? (you tell me)

OpenStudy (anonymous):

d=-17

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

So can you find the terms of the sequence ?

OpenStudy (solomonzelman):

When you go one term back, you add 17, when you go one term forward you subtract 17. OKAY ?

OpenStudy (oleg3321):

help me next solomon

OpenStudy (anonymous):

....:P okay i guess

OpenStudy (anonymous):

well, thank you for your help :)

OpenStudy (solomonzelman):

Type your sequence after you get it, so that I can check it. alright ?

OpenStudy (anonymous):

67,50,33,10,-1

OpenStudy (solomonzelman):

you got a little error, 67, 50, 33, \(\normalsize\color{red}{ 16}\), -1

OpenStudy (solomonzelman):

because 33 - 17 = 16 (and not 10) Good job though for the 1st 2 numbers, and good luck !

OpenStudy (anonymous):

oh whooops yea, i meant 17 :) lol

OpenStudy (solomonzelman):

you meant 16:)

OpenStudy (anonymous):

oh psht whatever it's all good now xP thankksss

OpenStudy (solomonzelman):

yeah, btw do you live in the very south somewhere or taking a summer school ?

OpenStudy (anonymous):

um why? i'm just taking this course online during the summer, yea

OpenStudy (solomonzelman):

I am just wondering.

OpenStudy (solomonzelman):

I'm not spying don't worry ;)